# Scheduling problems of aircraft on a same runway and dual runways

### v7

@@ VII-A Some necessary results
Proof: In this proof, we only discuss the statement (2) and statements (1)(3)(4) can be proved in a similar way. When $Tcf_{4}$ is a takeoff (landing) aircraft and $Tcf_{3}$ is relevant to $Tcf_{2}$, $\omega_{min}(\Phi_{a},Tcf_{5})=Y_{25}$. When $Tcf_{4}$ is a takeoff (landing) aircraft and $Tcf_{3}$ is relevant to $Tcf_{1}$, $T_{D}+D_{T}-Y_{13}\leq 0$ and $\omega_{min}(\Phi_{a},Tcf_{5})\geq Y_{25}+T_{D}+D_{T}-Y_{13}$. That is, $\omega_{min}(\Phi_{a},Tcf_{5})\geq Y_{25}+\min\{T_{D}+D_{T}-Y_{13},0\}$. Similarly, when $Tcf_{4}$ is a landing (takeoff) aircraft, it can be proved that $S_{3}(\Phi_{b})-S_{3}(\Phi_{a})\geq\min\{T_{D}+D_{T}-Y_{13},0\}+Y_{25}$ when $Tcf_{4}$ is relevant to $Tcf_{3}$ both in $\Phi_{a}$ and $\Phi_{b}$, and $S_{4}(\Phi_{b})-S_{4}(\Phi_{a})\geq\min\{T_{D}+D_{T}-Y_{13},T_{D}+D_{T}-Y_{24},0\}+Y_{25}$ when $Tcf_{4}$ is relevant to $Tcf_{2}$ in $\Phi_{a}$ and relevant to $Tcf_{5}$ or $Tcf_{3}$ in $\Phi_{b}$. That is, $\omega_{min}(\Phi_{a},Tcf_{5})\geq\min\{T_{D}+D_{T}-Y_{13},T_{D}+D_{T}-Y_{24},0\}+Y_{25}$.
@@
This theorem gives lower bounds for the time increments of the whole sequence when an aircraft is inserted, which can be used in the algorithm proposed later. Moreover, when defining $\omega_{min}$, since the aircraft $Tcf_{3}$ and $Tcf_{4}$ both have the possibilities of having no path from the last aircraft to themselves, we adopt the smaller value between $S_{3}(\Phi_{b})-S_{3}(\Phi_{a})$ and $S_{4}(\Phi_{b})-S_{4}(\Phi_{a})$ as the value of $\omega_{min}$ so as to implement a full space search for Algorithm 1 proposed later.
@@ VII-B Algorithm 1
$(i-3).$ According to the elements of the set $F_{inc}$, adjust the aircraft orders in the sequence $\phi_{(i-1)k}$ and insert the aircraft $Tcf_{1i}$ to generate a new sequence $\bar{\phi}_{ik}$ under the condition $F(\bar{\phi}_{ik})<f^{i}_{inc}$ such that $\bar{\phi}_{ik}$ contains the subsequence $\langle Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{1i},Tcf_{h_{3}},Tcf_{h_{4}}\rangle$ and $\langle h_{1},h_{2},h_{3},h_{4}\rangle\in F_{inc}$. Minimize $F(\bar{\phi}_{ik})$ based on the obtained theoretical results. Compare the values of all possible $F(\bar{\phi}_{ik})$ so as to find the optimal sequence of aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1i}$.
@@
If $f^{i}_{inc}-F(\phi_{i-1})=T_{0}$, by simple calculations, the inequality (21) holds when $(cl_{p},cl_{1i},cl_{j})=(2,5,4),(3,6,5),(4,6,5)$. Clearly, when $(cl_{p},cl_{1i},cl_{j})=(2,5,4),(3,6,5),(4,6,5)$, each aircraft $Tcf_{p}$ is a breakpoint aircraft and the aircraft $Tcf_{1i}$ is needed to be inserted to be between aircraft $Tcf_{p}$ and $Tcf_{j}$. To generate a subsequence $\langle Tcf_{p},Tcf_{1i},Tcf_{j}\rangle$, we can focus on the breakpoint aircraft and their trailing aircraft, and based on the theoretical results in previous sections, by adjusting the aircraft orders properly, the optimal sequence can be found under the condition that $F(\bar{\phi}_{ik})<f^{i}_{inc}$ in an easy way. If $f^{i}_{inc}-F(\phi_{i-1})=T_{0}+\delta$, the generation of a subsequence $\langle Tcf_{p_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ with $\langle{p_{1}},{j_{1}}\rangle\in F_{inc}$ would result in a new breakpoint aircraft or changes of the separation times between existing breakpoint aircraft and their trailing aircraft which might increase the value of the objective function $F(\cdot)$. Note for the separation times between breakpoint aircraft and their trailing aircraft that $T_{45}=T_{0}+\delta$ and $T_{ij}\geq 1.5T_{0}$ for all other $1\leq i<j\leq 6$. We can generate the whole sequence based on the subsequence $\langle Tcf_{p_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ according to time increments of the new breakpoint aircraft or changes of the separation times between breakpoint aircraft and their trailing aircraft under the condition that $F(\bar{\phi}_{ik})<f^{i}_{inc}$, and the optimal sequence can also be found in an easy way. Besides these two cases, the case of $f^{i}_{inc}-F(\phi_{i-1})>T_{0}+\delta$ as well as the takeoff sequences and mixed landing and takeoff sequences on a same runway can be discussed in a similar easy way.
From the above illustrations, we can first move the aircraft $Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}$ to proper places by equivalent transformations including breakpoint-drift equivalent transformations and then use the extraction operation and insertion operation to generate the sequence $\bar{\phi}_{ik}$ with a focus on the breakpoint aircraft, their trailing aircraft and the landing-takeoff/takeoff-landing transitions as well as the properties of class-monotonically-decreasing sequences on the basis of the optimal sequence.
@@
Since each $\phi_{(i-1)k}$ is an optimal sequence for aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1(i-1)}$, if there is no utilizable subsequence $\langle Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}\rangle$ formed by four consecutive aircraft of $Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}$ in all $\phi_{(i-1)k}$, the generation of $\langle Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}\rangle$ in $\phi_{(i-1)k}$ would increase the value of the objective function $F(\cdot)$, which is the reason why only the aircraft combinations satisfying (20) are taken into account. Moreover, the optimal insertion of $Tcf_{1(i-1)}$ might usually generate a new breakpoint/landing-takeoff/takeoff-landing transition or change an existing separation time between a breakpoint aircraft and its trailing aircraft with some necessary aircraft order adjustments. These properties and related discussions in previous sections can be used to reduce the computation amount of Algorithm 1.
@@
Based on the above analysis, we can discuss the optimality of the sequence $\bar{\phi}_{ik}$ in the following cases.
@@
Scenario 1. Landing/takeoff sequences. Let $Nb_{\bar{\phi}_{ik}}$ and $Nb_{\phi_{(i-1)k}}$ denotes the numbers of breakpoint aircraft in the sequences $\bar{\phi}_{ik}$ and $\phi_{(i-1)k}$.
@@
Scenario 2. Mixed landing and takeoff sequences on a same runway. Let $Nt_{\bar{\phi}_{ik}}$ and $Nt_{\phi_{(i-1)k}}$ denotes the numbers of landing-takeoff and takeoff-landing transitions in the sequences $\bar{\phi}_{ik}$. and $\phi_{(i-1)k}$.
@@
Since different classes might correspond to different time increments, we can search for the desired aircraft of proper classes in $F_{inc}$ in ascending order of time increments to minimize the value of the objective function $F(\cdot)$. Moreover, note that the separation times between aircraft of different classes are time-invariant and the number of the total aircraft classes is not very large. The computation cost of time increments is also not very large and can even be calculated out offline.
@@
Remark 37
Due to the constraints of time windows of aircraft, breakpoint aircraft and resident-point aircraft might be generated which might increase the value of the objective function $F(\cdot)$. To analyze the optimal sequence of the aircraft, the emphasis should be imposed on the breakpoint aircraft and the resident-point aircraft as well as the separation times between the consecutive aircraft of the same class that are larger than $T_{0}$ by trying not to increase their numbers.
@@ VII-C Algorithm 2
Step 2. (2.1) According to the landing/takeoff time increments of block subsequences, classify all possible block subsequences into several sets, which can be processed offline.
When our algorithm is applied, in the landing/takeoff sequence of each block subsequence, we consider only the case of one breakpoint aircraft and the case of two or more breakpoint aircraft in a block subsequence can be dealt with by combing the operations of extraction and insertion to decrease the value of the objective functions based on the analysis of the case of one breakpoint aircraft for block subsequences. Or alternatively, the case of two or more breakpoint aircraft in a block subsequence can be dealt with offline by using the same method for the case of one breakpoint aircraft.
As discussed in Sec. VI, if the number of breakpoint aircraft in a block subsequence is fixed, the number of all possible block subsequences of different class combinations is not large according to the minimum separation time standards at Heathrow Airport and for the RECAT-EU system, when resident-point aircraft are not taken into account.
(2.2) Focus on combinations of different block subsequences, and study the landing/takeoff time increments of any two given consecutive block subsequences, where the last two aircraft of the first block subsequence is the first two aircraft of the second block subsequence. According to the landing/takeoff time increments, classify all possible two consecutive subsequences into several sets.
In our simulations for the case of $F(\Phi_{a})\geq F(\Phi_{b})$, the combination of a D-block subsequence and a T-block subsequence is commonly used, which has the form of $\langle DB,TB\rangle$ where $DB$ denotes a D-block subsequence and $TB$ denotes a T-block subsequence, the first two aircraft of $DB$ have the same landing/takeoff time, and the last two aircraft of $TB$ have the same landing/takeoff time (see Tables III and I for related discussions). The landing time increment of D-block subsequence in such a combination is a key quantity that needs to be studied for the optimal sequence.
Step 3. (3.1) Match the takeoff aircraft with each landing sequence in $\Lambda_{a}$ to generate a new sequence, denoted by $\Phi_{c}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n+m}\rangle$, and optimize $\Phi_{c}$ such that $F(\Phi^{c}_{LA},Sr_{LA}(\Phi_{c}))=F(\Phi_{a})$, where $\Phi^{c}_{LA}$ and $Sr_{LA}(\Phi_{c})$ denote the landing sequence and the corresponding landing time vector in $\Phi_{c}$.
@@
Here, it should be noted that each sequence of $\Lambda_{a}$ is required to be considered. Actually, we usually need to consider only the impact of the local differences of the sequence from other sequences instead of the whole sequence, which might not need a lot of the computation amount.
(3.2) Let $n_{z}$ denote the number of the aircraft of $\Phi_{c}$ which take off in $(t_{0}+F(\Phi_{a}),+\infty)$. Insert the last $n_{z}$ aircraft of $\Phi_{c}$ into the subsequence of $\Phi_{c}$, $\Phi^{sub}_{c}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n+m-n_{z}}\rangle$, to generate a new sequence $\Phi_{c3}$ and minimize $F(\Phi_{c3})$ without changing the orders of the landing aircraft based on the sets defined Step 2 by combing the length, the initial redundant time and the landing/takeoff time increment of each block subsequence. Note here that the sequence $\Phi_{c3}$ can contain D-block subsequences.
The optimal value of the objective function $F(\cdot)$ for the optimization problem (6) lies in $[F(\Phi_{a}),F(\Phi_{c3})]$.
Step 4. The main idea of this step is to implement a full space search for an optimal solution $(\Phi_{c4},Sr(\Phi_{c4}))$ for the optimization problem (6) under the condition that $F(\Phi_{a})\leq F(\Phi_{c4})\leq F(\Phi_{c3})$ starting from the initial condition that $\Phi^{c4}_{LA}\in\Lambda_{a}$. Note that the set $\Lambda_{a}$ used here does not contain the information of the landing/takeoff times.
Since different D-block subsequences might have different landing time increments, we can search for the optimal solutions for the optimization problem (6) by increasing the sum of the landing time increments of D-block subsequences of the sequence $\Phi_{c4}$. Based on this idea, since $F(\Phi_{a})\leq F(\Phi^{c4}_{LA})\leq F(\Phi_{c4})$, the condition $F(\Phi_{a})\leq F(\Phi_{c4})\leq F(\Phi_{c3})$ can be relaxed to be $F(\Phi_{a})\leq F(\Phi^{c4}_{LA})\leq F(\Phi_{c3})$.
If the sum of the landing time increments of D-block subsequences of $\Phi_{c4}$ is larger than $F(\Phi_{c3})-F(\Phi_{a})$, then it follows that $F(\Phi_{c4})>F(\Phi_{c3})$, implying that $(\Phi_{c4},Sr(\Phi_{c4}))$ is not an optimal solution for the optimization problem (6).
@@
Different from the approach described above, there is another search approach to generate the optimal sequence directly according to block subsequences and the combinations of block subsequences discussed in Step 2, which might have more computation amount.
@@
The main idea of Algorithm 2 is to decompose the aircraft sequence into several block subsequences and fully explore combinations of the block subsequences starting from an optimal landing/takeoff subsequence to consider the optimization problem (6). Note that the number of all possible block subsequences of different class combinations is not large according to the minimum separation time standards at Heathrow Airport and for the RECAT-EU system, when resident-point aircraft are not taken into account. The use of block subsequences might significantly reduce the computation amount of the algorithm.

### v8

@@ VII-A Some necessary results
Proof: In this proof, we only discuss the statement (2) and statements (1)(3)(4) can be proved in a similar way. When $Tcf_{4}$ is a takeoff (landing) aircraft and $Tcf_{3}$ is relevant to $Tcf_{2}$ in $\Phi_{a}$, $\omega_{min}(\Phi_{a},Tcf_{5})=Y_{25}$. When $Tcf_{4}$ is a takeoff (landing) aircraft and $Tcf_{3}$ is relevant to $Tcf_{1}$ in $\Phi_{a}$, $T_{D}+D_{T}-Y_{13}\leq 0$ and $\omega_{min}(\Phi_{a},Tcf_{5})\geq Y_{25}+T_{D}+D_{T}-Y_{13}$. That is, $\omega_{min}(\Phi_{a},Tcf_{5})\geq Y_{25}+\min\{T_{D}+D_{T}-Y_{13},0\}$. Similarly, when $Tcf_{4}$ is a landing (takeoff) aircraft, it can be proved that $S_{3}(\Phi_{b})-S_{3}(\Phi_{a})\geq\min\{T_{D}+D_{T}-Y_{13},0\}+Y_{25}$ when $Tcf_{4}$ is relevant to $Tcf_{3}$ both in $\Phi_{a}$ and $\Phi_{b}$, and $S_{4}(\Phi_{b})-S_{4}(\Phi_{a})\geq\min\{T_{D}+D_{T}-Y_{13},T_{D}+D_{T}-Y_{24},0\}+Y_{25}$ when $Tcf_{4}$ is relevant to $Tcf_{2}$ in $\Phi_{a}$ and relevant to $Tcf_{5}$ or $Tcf_{3}$ in $\Phi_{b}$. That is, $\omega_{min}(\Phi_{a},Tcf_{5})\geq\min\{T_{D}+D_{T}-Y_{13},T_{D}+D_{T}-Y_{24},0\}+Y_{25}$.
@@
This theorem gives lower bounds for the time increments of the whole sequence when an aircraft is inserted, which can be used in Algorithm 1. Moreover, when defining $\omega_{min}$, since the aircraft $Tcf_{3}$ and $Tcf_{4}$ both have the possibilities of having no path from the last aircraft to themselves, we adopt the smaller value between $S_{3}(\Phi_{b})-S_{3}(\Phi_{a})$ and $S_{4}(\Phi_{b})-S_{4}(\Phi_{a})$ as the value of $\omega_{min}$ so as to implement a full space search for Algorithm 1.
@@ VII-B Algorithm 1
$(i-3).$ According to the elements of the set $F_{inc}$, adjust the aircraft orders in the sequence $\phi_{(i-1)k}$ and insert the aircraft $Tcf_{1i}$ to generate a new sequence $\bar{\phi}_{ik}$ under the condition $F(\bar{\phi}_{ik})<f^{i}_{inc}$ such that $\bar{\phi}_{ik}$ contains the subsequence $\langle Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{1i},Tcf_{h_{3}},Tcf_{h_{4}}\rangle$ and $\langle h_{1},h_{2},h_{3},h_{4}\rangle\in F_{inc}$. Minimize $F(\bar{\phi}_{ik})$ based on the obtained theoretical results, and compare the values of all possible $F(\bar{\phi}_{ik})$ so as to find the optimal sequence of aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1i}$.
@@
If $f^{i}_{inc}-F(\phi_{i-1})=T_{0}$, by simple calculations, the inequality (21) holds when $(cl_{p},cl_{1i},cl_{j})=(2,5,4),(3,6,5),(4,6,5)$. Clearly, when $(cl_{p},cl_{1i},cl_{j})=(2,5,4),(3,6,5),(4,6,5)$, each aircraft $Tcf_{p}$ is a breakpoint aircraft and the aircraft $Tcf_{1i}$ is needed to be inserted to be between aircraft $Tcf_{p}$ and $Tcf_{j}$. To generate a subsequence $\langle Tcf_{p},Tcf_{1i},Tcf_{j}\rangle$, we can focus on the breakpoint aircraft and their trailing aircraft, and based on the theoretical results in previous sections, by adjusting the aircraft orders properly, the optimal sequence can be found under the condition that $F(\bar{\phi}_{ik})<f^{i}_{inc}$ in an easy way. If $f^{i}_{inc}-F(\phi_{i-1})=T_{0}+\delta$, the generation of a subsequence $\langle Tcf_{p_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ with $\langle{p_{1}},{j_{1}}\rangle\in F_{inc}$ would result in a new breakpoint aircraft or some changes of the separation times between existing breakpoint aircraft and their trailing aircraft. Considering the values of the separation times between all possible breakpoint aircraft and their trailing aircraft, note that $T_{45}=T_{0}+\delta$ and $T_{ij}\geq 1.5T_{0}$ for all $1\leq i<j\leq 6$ with $(i,j)\neq(4,5)$. We can generate the whole sequence based on the subsequence $\langle Tcf_{p_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ by considering the total time increments caused by the insertion of the aircraft $Tcf_{1i}$ under the condition that $F(\bar{\phi}_{ik})<f^{i}_{inc}$.
@@
Since each $\phi_{(i-1)k}$ is an optimal sequence for aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1(i-1)}$, if there is no utilizable subsequence $\langle Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}\rangle$ formed by four consecutive aircraft of $Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}$ in all $\phi_{(i-1)k}$, the generation of $\langle Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}\rangle$ in $\phi_{(i-1)k}$ would increase the value of the objective function $F(\cdot)$ and $F(\bar{\phi}_{ik})\geq f^{i}_{inc}$ when $\omega_{min}(\langle Tcf_{h_{1}},Tcf_{h_{2}},Tcf_{h_{3}},Tcf_{h_{4}}\rangle,Tcf_{1i})\geq f^{i}_{inc}-F(\phi_{i-1})$. This is the reason why only the aircraft combinations satisfying (20) are taken into account. Moreover, the optimal insertion of $Tcf_{1(i-1)}$ might usually generate a new breakpoint/landing-takeoff/takeoff-landing transition or change an existing separation time between a breakpoint aircraft and its trailing aircraft with some necessary aircraft order adjustments. These properties and related discussions in previous sections can be used to reduce the computation amount of Algorithm 1.
@@
Based on the above analysis, we can combine the following cases to discuss the optimality of the sequence $\bar{\phi}_{ik}$.
@@
Scenario 1. Landing/takeoff sequences. Let $Nb_{\bar{\phi}_{ik}}$ and $Nb_{\phi_{(i-1)k}}$ denote the numbers of breakpoint aircraft in the sequences $\bar{\phi}_{ik}$ and $\phi_{(i-1)k}$.
@@
Scenario 2. Mixed landing and takeoff sequences on a same runway. Let $Nt_{\bar{\phi}_{ik}}$ and $Nt_{\phi_{(i-1)k}}$ denote the numbers of landing-takeoff and takeoff-landing transitions in the sequences $\bar{\phi}_{ik}$. and $\phi_{(i-1)k}$.
@@
Since different class combinations might correspond to different time increments, we can search for the desired aircraft of proper classes in $F_{inc}$ in ascending order of time increments to minimize the value of the objective function $F(\cdot)$. Moreover, note that the separation times between aircraft of different classes are time-invariant and the number of the total aircraft classes is not very large. The computation cost of time increments is also not very large, which can be calculated out offline.
@@ VII-C Algorithm 2
Step 2. (2.1) According to the landing/takeoff time increments of block subsequences and the number of breakpoint aircraft, classify all possible block subsequences into several sets, which can be processed offline.
When the number of breakpoint aircraft in a block subsequence is fixed, all possible block subsequences can be obtained by an enumeration method. It should be noted that the block subsequences with more breakpoint aircraft can be generated based on the block subsequences with less breakpoint aircraft. It should be also noted that the number of all possible block subsequences of different class combinations and the range of time lengths and the landing/takeoff increments of block subsequences are both not large under the minimum separation time standards at Heathrow Airport and for the RECAT-EU system are considered without considering resident-point aircraft, when the number of breakpoint aircraft in a block subsequence is fixed.
(2.2) Focus on combinations of different block subsequences, and study the landing/takeoff time increments of any two given consecutive block subsequences, where the last two aircraft of the first block subsequence is the first two aircraft of the second block subsequence. According to the landing/takeoff time increments and the number of breakpoint aircraft, classify all possible two consecutive subsequences into several sets.
In our simulations for the case of $F(\Phi_{a})\geq F(\Phi_{b})$, the combination of a D-block subsequence and a T-block subsequence is frequently used, which has the form of $\langle DB,TB\rangle$ where $DB$ denotes a D-block subsequence and $TB$ denotes a T-block subsequence, the first two aircraft of $DB$ have the same landing/takeoff time, and the last two aircraft of $TB$ have the same landing/takeoff time (see Tables III and I for related discussions). The time length and the landing time increment of D-block subsequence in such a combination are key quantities that needs to be studied for the optimal sequence.
Step 3. (3.1) Match the takeoff aircraft with a landing sequence in $\Lambda_{a}$ to generate a new sequence, denoted by $\Phi_{c}=\langle Tcf_{c1},Tcf_{c2},\cdots,Tcf_{c(n+m)}\rangle$, and minimize $\Phi_{c}$ under the condition that $F(\Phi^{c}_{LA},Sr_{LA}(\Phi_{c}))=F(\Phi_{a})$, where $\Phi^{c}_{LA}$ and $Sr_{LA}(\Phi_{c})$ denote the landing sequence and the corresponding landing time vector in $\Phi_{c}$.
@@
(3.2) Let $n_{z}$ denote the number of the aircraft of $\Phi_{c}$ which take off in $(t_{0}+F(\Phi_{a}),+\infty)$. Insert the last $n_{z}$ aircraft of $\Phi_{c}$ into the subsequence of $\Phi_{c}$, $\Phi^{sub}_{c}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n+m-n_{z}}\rangle$, to generate a new sequence $\Phi_{c3}$ and minimize $F(\Phi_{c3})$ without changing the orders of the landing aircraft based on the sets defined Step 2 by combing the capacity, the time length, the initial redundant time and the landing/takeoff time increment of each block subsequence. Note here that the sequence $\Phi_{c3}$ can contain D-block subsequences.
It should be noted that the minimum value of the objective function $F(\cdot)$ for the optimization problem (6) lies in $[F(\Phi_{a}),F(\Phi_{c3})]$ and the final optimal sequence usually contains at least $n_{z}-1$ D-block subsequences, which can be used as an initial condition to find a feasible solution near the optimal solutions to narrow the search range in step 4.
Step 4. The main idea of this step is to implement a full space search for an optimal solution $(\Phi_{c4},Sr(\Phi_{c4}))$ for the optimization problem (6) under the condition that $F(\Phi_{a})\leq F(\Phi_{c4})\leq F(\Phi_{c3})$ starting from the initial condition that $\Phi^{c4}_{LA}\in\Lambda_{a}$ base on T-block and D-block subsequences.
Since different D-block subsequences might have different landing time increments, we can search for the optimal solutions for the optimization problem (6) by increasing or decreasing the number and the landing time increments of D-block subsequences of the sequence $\Phi_{c4}$. Based on this idea, since $F(\Phi_{a})\leq F(\Phi^{c4}_{LA})\leq F(\Phi_{c4})$, the condition $F(\Phi_{a})\leq F(\Phi_{c4})\leq F(\Phi_{c3})$ can be relaxed to be $F(\Phi_{a})\leq F(\Phi^{c4}_{LA})\leq F(\Phi_{c3})$.
When optimizing the sequence, we can first consider the block subsequences containing the breakpoint aircraft since the number of all possible block subsequences containing breakpoint aircraft is usually larger than the number of all possible block subsequences containing no breakpoint aircraft without considering the aircraft classes.
@@
There are several method, e.g., binary search method, to realize the search in this step, and the most basic method is to generate the optimal sequence completely according to block subsequences and the combinations of block subsequences discussed in Step 2, which might have more computation amount. To reduce the computation amount, we can use the number of breakpoint aircraft to partition the search space as stated in Algorithm 1.
@@
The main idea of Algorithm 2 is to decompose the aircraft sequence into several block subsequences and fully explore combinations of the block subsequences starting from an optimal landing/takeoff subsequence to consider the optimization problem (6). Note that the number of all possible block subsequences of different class combinations is usually not large under the minimum separation time standards at Heathrow Airport and for the RECAT-EU system without considering resident-point aircraft. The use of block subsequences might significantly reduce the computation amount of the algorithm.